change in Ph in buffers

part A

What is the pH of a buffer prepared by adding 0.607mol of the weak acid HA to 0.608mol of NaA in 2.00 L of solution? The dissociation constant Ka of HA is 5.66
Answer #1

a).pH = pKa + log(A-/HA)

pKa = -logKa = 6.247

pH = 6.247+log(0.608/.607) = 6.2477


b) after addition of 0.150 mol HCl, A- decreases from 0.608 mol to 0.458 mol per 2L; the HA increases from 0.607 mol to 0.757 mol
pH = pKa + log(0.458/0.757) = pKa - 0.22

pH = 6.027
c) A- increases from 0.608 mol to 0.803 mol; HA decreases from 0.607 mol to 0.412 mol.
pH = pKa + log(0.803/0.412) = pKa + 0.29

pH = 6.537
Answer #2

part A

What is the pH of a buffer prepared by adding 0.607mol of the weak acid HA to 0.608mol of NaA in 2.00 L of solution? The dissociation constant Ka of HA is 5.66